OpenStudy (anonymous):

I have a problem that looks like the following and I need to determine y for 0 and n for 0. can you help me? n y 2 8 7 3 5 5 0 0 y =

OpenStudy (amistre64):

2-7 = -5 8-3 = 5 ; so far that might indicate a slope of -1 5-2 = 3 5-8 =-3 ; its still a -1 ratio i believe your sequence can be modeled after a linear equation: y = -n + C 5 5 5 = -5 + 10 y = -n+10

OpenStudy (amistre64):

when n=0; y=10 when y=0; n=10